OPRE 3360 - Managerial Methods in Decision Making Under Uncertainty Chapter 2- Frequency, Relative Frequency and Percent Frequency Distributions 3. A partial frequency sample are given. Calculate the missing data on both tables distribution and a partial cunulative frequency distribution for a particular Class 1 -20 21 40 41 60 61-8013 81 100 Frequency Relative Frequency 25 Cumulative Frequency Cumulative Relative Frequency Class S 20 s 40 s 60 s 80 100 17 175 1.00
Answer #1

Computational Table:

Class Frequency Relative Frequency Cumulative frequency Cumulative relative frequency
1 - 20 3 3/75 = 0.04 3 0.04
21 - 40 17 - 3 = 14    14/75 = 0.19 17 0.04+0.19 = 0.23
41 - 60 25 25/75 = 0.33 17+25 = 42 0.33+0.23 = 0.56
61 - 80 13 13/75 = 0.17 42+13 = 55 0.56 + 0.17 = 0.73
81 - 100 20 20/75 = 0.27 55+20 = 75 0.73+0.27 = 1.00
Total 75 1.00

Cumulative frequency for 81-100 = 75, That is, Total Frequency = 75

Cumulative frequency for 21-40 = frequency of 1-20 class + frequency of 21-40 class

17 = 3 + frequency of 21-40 class

Therefore, frequency of 21-40 class = 17-3 = 14

frequency of 81-100 class = Total - All classes

frequency of 81-100 class = 75 - (3 + 14 + 25+13) = 20